I think of a pair of numbers. If i add 11 to the first, ibobtain a number which is twice the second. If i add 20 to the second, i obtain a number which is twice the first. What are the numbers?
Answer by ankor@dixie-net.com(22740)

I think of a pair of numbers.
a & b
If i add 11 to the first, i obtain a number which is twice the second.
a + 11 = 2b
a = 2b - 11
If i add 20 to the second, i obtain a number which is twice the first.
b + 20 = 2a
Replace a with (2b-11)
b + 20 = 2(2b-11)
b + 20 = 4b - 22
20 + 22 = 4b - b
42 = 3b
b = 42/3
b = 14
I'll let you find a, check your solutions in both statements

What are the numbers?